From: Mon 1/26/09


  • UofC ATLAS meeting (3rd HEP)
  • FDR meeting (phone EsNet: call 1-510-883-7860, code 888726#; esnet video, ad hoc 888726)
  • MWT2 meeting (phone in Robert's office)


  • ATLAS ADCoS (prodsys shifters) meeting (phone CERN, Xavier Espinal) - report, shifter on duty
  • ATLAS TAG developers meeting (phone CERN, David Malon) - late, overlapping
  • ATLAS shift
    • group ticket for Tier2?
  • provide updated ATLAS-wn to Xin and coordinate update
  • troubleshooting:
    • "RE: set OU_OSCER_ATLAS to test, KV and release/RHEL comatibility "


  • ATLAS DB meeting (phone CERN) (skipped)
  • ATLAS facilities meeting (phone EsNet)
  • problem solving
    • Anton K: getting files on WN from Tier3 SE (PFN problems on edge5, forgot that /pfn/ is not working for gridftp server)
    • Esteban O: Athena failing because of missing file returning GUID: lookup in Panda monitor, belogns to a Ghost dataset (only at BNL and no files in it), but the file is on disk (retrievable querying LFC)
  • Started report on 1/12 analysis jobs
  • Site surveillance during ATLAS-wn update



-- MarcoMambelli - 29 Jan 2009
Topic revision: r3 - 31 Jan 2009, MarcoMambelli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback